wiki:UC20200707

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 07/07/2020

Attending: Mark Klaene (APO), Jamey Eriksen (APO), John Bally (CU), Moire Prescott (NMSU), Eric Bellm (UW), Kevin Schlaufman (JHU), Anne Verbiscer (UVa), Sarah Tuttle (UW), Ben Williams (UW), Misty Bentz (GSU), Kal Kadlec (UW), Derek Buzasi (FGCU), Chip Kobulnicky (UWy), Denise Stephens (BYU), Russet McMillan (APO)


We started by introducing Denise Stephens, the Users Committee representative from Brigham Young University (BYU). Although BYU's membership in ARC doesn't officially take effect until January 2021, Denise will be participating in the UC meetings between now and then in an effort to get up to speed on various aspects of the 3.5m operations. We are delighted to welcome BYU as our newest ARC member - welcome Denise!

User feedback and comments from institutional representatives:

  • Oklahoma - nothing to report (by email)
  • NMSU - nothing to report
  • Colorado - nothing to report
  • Wyoming - nothing to report
  • Seattle - no report
  • Washington - nothing to report
  • NAPG - no report
  • JHU - nothing to report
  • UVa - nothing to report
  • Georgia State - nothing to report
  • FGCU - nothing to report

Telescope and Instruments Report:

Jamey's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 5/28/2020 - 7/03/2020

0) Overview

APO has been open for night operations since June 1. We are still restricting all visitors to the site. Following the state of New Mexico and NMSU guidelines, we also remain under travel restrictions ourselves. On-site staffing levels have been increased slightly from the minimal staff allowed on site before 6/1/20. Currently, as NMSU is in Phase 3 of the Return to Research, APO is allowed between 35-50% of staffing. The NMSU numbers, coupled with the state of New Mexico restrictions for a non-essential classified business like APO, which allow staffing levels of 50% of the Fire Marshall determined maximum occupancy of a building, has allowed us to have a more people on site at any given time. Weather for the month of June has been quite good as we have yet to be affected by onset of the monsoon rains. We did not lose some nights to precipitation or high humidity. The early part of the month experienced several nights of elevated tree pollen. We are experiencing an above average year of Miller moths and early June saw an abnormal peak in the ladybug population.

1) Telescope

The telescope has been performing as expected. There was a recent issue with the NA2 rotator but we believe that an adjustment to the preload fixed the issue. Telescope summer shutdown will begin on 7/6/20 with a return to science expected on 7/15/20 (weather dependent). There will be no recoating of any of the mirrors during shutdown this year.

2) Instruments

DIS: working as expected. Scattered light is still an issue with the instrument. Turret (and/or detent) motor appears to be acting up, requiring more than one home resets to clear the issue.

TripleSpec: powered on, cold and operational.

Agile: powered on, cold and operational.

Echelle: powered on, cold and operational. The inter-order light (IOL) for red and blue continue to drop, but at a slower rate than the previous month. See link provided by Bill K. for plots of the IOL over time: http://users.apo.nmsu.edu/~bketzeba/instruments/echelle/echelle_IOL_update.html

NIC-FPS: the ion pump was replaced and the instrument is cold and operational.

ARCTIC: powered on, cold, and operational.

3) ARCSAT

ARCSAT is operational and has been used for observations this past month. A low flux level streaking issue in SurveyCam images is being investigated.


Additional telescope and instrument discussion:

We have been back in operation since June 1, adhering to Governor of NM guidelines and NMSU's Return to Research guidelines. These are slightly in conflict but we worked out the maximum occupancy numbers for each building, and 50% of max occupancy is currently allowed. The weather has been quite good for June. Monsoon rains seem to be starting, and we are seeing extra moths and ladybugs at the site. The 3.5m telescope is performing as expected, and shutdown tasks are going well. The primary mirror was cleaned, and collimation and reassembly are on track. We are on schedule for our planned return to science on 7/15, weather permitting. As a reminder, the mirrors will not be recoated this summer due to the KPNO shutdown, but we will try to get in line for next summer as this will be our sixth year since our last recoating.

Instrument updates beyond what is covered in the monthly report above pertained to DIS. Bill will look at the DIS turret (or detent) motor issue after summer shutdown. Russet reported that the scattered light in DIS-B is improving; the right side of is DIS-B better to the point that users can place lines there. DIS-R is still deteriorating as usual, but it remains better than DIS-B. Russet and/or Bill will to look into how much one needs to move target lines on the right side of DIS-B. Misty asked what is the timescale for variations in scattered light? Russet replied that it evolves more in warm weather and it improves or stabilizes in wintertime. Thus, having DIS-B improve during summer is confusing, while seeing DIS-R gradually deteriorate is as expected. If we pump on it (planned for late Sept/Oct), we would expect to see rapid changes - on the timescale of hours - right after pumping, and then slowing to days and stabilizing. We will pump on it after the hot weather is over. We are currently checking it once a week, sometimes seeing a little evolution, but in general we are seeing more changes during warmer weather.


APO and the COVID-19 virus

Current staffing restrictions haven’t changed much. The NM Governor’s order on July 1 now requires a 14-day self-quarantine for anyone traveling into NM from out of state (both residents and non-residents), which has an impact on both the SDSS and 3.5m observing staff. We are going to be operating short-staffed (in terms of 3.5m obs-specs) with no margins for the near future (unclear for how long). This means that the obs-specs are working longer hours and more days per shift. We are actively discussing both short- and long-term solutions to this problem. For example, we have volunteers from the SDSS 2.5m obs-specs to be the eyes and ears if we don’t have an on-site 3.5m operator, e.g., we could have a remote 3.5m obs-spec with a SDSS obs-spec as the on-site person.

UC reps should make users aware that there is a chance that observing on any given night may not happen due to this situation. We may also put a statement to this effect in the auto-generated reminder emails. Please be patient with and kind to your observing specialists.


Q3 3.5m scheduling:

Users can request remaining OPEN/DD time in Q3 by sending request and justifications to Nancy, Ben, Russet. There isn’t much remaining (Aug 16A, Sep 29A), although there is a handful of 1-2-hour blocks in late July and early Aug.

UC reps: notify users that they should let Ben, Russet and Nancy know if they are interested in requesting any of the remaining OPEN/DD time. For an existing program this requires simply sending an email; if the request is not associated with an existing program already scheduled in Q3 they will ask for a short science justification.


Q3 ARCSAT scheduling:

The Q3 schedule is posted, and there are several weeks of OPEN time remaining (July 13-19, Jul 27-Aug 2, Sep 21-27). Bill plans to get the Shelyak echelle on ARCSAT hopefully by the end of the quarter. He will be working on it in August and the first week of September is dedicated to engineering. Russet reported that there is an error in the ARCSAT schedule, with all OPEN weeks currently incorrectly listed as SurveyCam. She will double-check which camera will be available when.

UC reps: have your users let Ben, Nancy and Russet know if they are interested in requesting any of the remaining ARCSAT time.


KOSMOS Update:

Sarah reported that her group has been allowed back in the lab as of late last week. They are working to get the new KOSMOS dewar assembled. They will then test things will the new slits. The amount of time they can spend in the lab is still limited; her group is still mostly teleworking, and scheduling lab time as necessary. They are still trying to address the vacuum problem, and are ramping back up with a list of tasks to understand the problem.


ARCTIC Diffuser Rotation

For the diffuser rotation mechanism, the parts are in the shop but nothing is being assembled yet but hopefully soon. This will be done in the same lab as the KOSMOS work, but the lab is large enough and well-ventilated enough that it should be okay. Kal is waiting to make the call on when to work on it as stuff comes in. Ben asked if there is a timeline for when the diffuser rotator will be ready (e.g., ready for use in Q1 2021?). Kal replied that the part that will take time is getting the motor running and programmed. It's difficult to estimate a commissioning date given the limited lab time.


Status Update on Near-Term Action Items from Strategic Planning Meeting

  • DAWG: There is no update this month. Nancy is trying to reactivate the group.
  • Getting narrow-band filters for ARCTIC: Nancy received an updated quote from the vendor and expects to order the filters this week. (She is also getting a quote for an additional request from John Bally). She has unofficial price quotes and is in contact with Rene Walterbos and John Bally on strategy. We have to make a choice between two possible bandpasses, 30A vs. 50-75A band, with 30A band being significantly more expensive. This comes down to whether one cares a lot about isolating Ha vs. [NII] vs. don’t care. If users have strong feelings on this, they should let Nancy know. Otherwise, we will likely go with the wider 75A Halpha filter. John Bally asked when the existing UNM Halpha filters were made? He doesn't know the provenance of those filters, but they are quite nice, and would like to find out the vendor. The price quotes Nancy received seemed a bit high, although he realizes that the number of vendors has shrunk. Ben asked if we go very narrow for Ha would we also want to go very narrow for OIII? John replied yes, because we want to isolate the continuum. This is a unique opportunity with this f/10 telescope, we can go narrower. Russet pointed out that part of the problem with going very narrow is that it is great for Galactic work but rules out extragalactic sources with even a slight redshift. Then we will end up with a huge library of Ha filters.

UC reps should instruct their users that if they have strong feelings about whether we should go with very narrow or less narrow Ha etc. filters for ARCTIC, they should notify Nancy ASAP.

  • Science Communications Interns: Nancy and Emily will be doing a pilot "science highlights hangout” with NMSU users (for eventual posting on YouTube and/or other social media). If this works, they will try it with other ARC institutions. They will also be doing some occasional short posts about different observing programs going on (with the PI’s permission first), so they will contact people shortly before their observing run to ask if they’d be willing to share a raw image or spectrum (or a candid shot of people observing, or screenshot of TUI).

ACTION ITEMS:

Open action items from previous meetings:

  • ARCTIC filter order - Nancy. Status: OPEN. Nancy will contact Rene about the UNM filters.

New action items from this meeting:

  • UC reps: notify users that we are operating with fewer obs-specs than normal and there is a chance this could impact our observing schedule. Also, please ask users to be patient with the night staff, and to respond to the reminder emails in a timely fashion.
  • UC reps: notify users that there is a limited amount of OPEN/DD time remaining on both the 3.5m and ARCSAT during Q3. If interested in requesting some time, please follow standard protocols for making such requests.
  • UC reps: instruct users that if they have strong feelings about whether we should go with very narrow or less narrow Ha etc. filters for ARCTIC, they should notify Nancy ASAP.
  • UC reps: remind all users to send their list of refereed/submitted publications from 2019-present to Nancy by 7/31/2020. (see below under All Other Business)

All Other Business

This is a reminder to all users that they should send Nancy a listing of all refereed publications and student theses and dissertations that made use of data acquired with the 3.5m and/or ARCSAT published from 2019 to the present (including “submitted”). She needs to receive this by July 31.

This was Mark Klaene's last APO Users Committee meeting prior to his retirement. He expressed his thanks to everyone for your help in improving the observatory, and said he learned a lot from everyone. [Note added in proof from Nancy: THANK YOU MARK for your participation in these Users Committee telecons over so many years! You have done a terrific job of representing the site perspective on these calls, and I know I speak for all of us when I express my deepest thanks for your dedication to enabling all kinds of cool science that our users have been able to accomplish with the 3.5m telescope and ARCSAT.]


Next meeting:

The next meeting will be on Monday August 3 at 10:30 am MDT.


Note: See TracWiki for help on using the wiki.